OpenStudy (anonymous):

How is 1 the coefficient of x^8y^6

Directrix (directrix):

One is the multiplicative identity and one times any quantity such as x^8y^6 = 1* x^8y^6 = x^8y^6. I that it that x^8y^6 is a term in a polynomial. Nothing changes in value.

Directrix (directrix):

coefficient [ˌkəʊɪˈfɪʃənt] n 1. (Mathematics) Maths a. a numerical or constant factor in an algebraic term the coefficient of the term 3xyz is 3 http://www.thefreedictionary.com/coefficient